### v1

# Werewolf: A Straightforward Game Framework with TTS for Improved User Engagement

Abstract
The growing popularity of social deduction game systems for both business applications and AI research has greatly benefited from the rapid advancements in Large Language Models (LLMs), which now demonstrate stronger reasoning and persuasion capabilities. Especially with the raise of DeepSeek R1 and V3 models, LLMs should enable a more engaging experience for human players in LLM-agent-based social deduction games like Werewolf. Previous works either fine-tuning, advanced prompting engineering, or additional experience pool to achieve engaging text-format Werewolf game experience. We propose a novel yet straightforward LLM-based Werewolf game system with tuned Text-to-Speech(TTS) models designed for enhanced compatibility with various LLM models, and improved user engagement. We argue with ever enhancing LLM reasoning, extra components will be unnecessary in the case of Werewolf.
Keywords Social Deduction Game Human-AI Interaction Human-Computer Interaction
1 Introduction
The Werewolf game, a widely recognized social deduction game, continues to attract increasing attention from AI researchers, particularly within the natural language processing (NLP) community. Prior work has introduced dedicated datasets for this domain [1] and leveraged them to fine-tune models for experience retrieval, enhancing large language model (LLM) performance in gameplay [2, 3]. These approaches often integrate external modules to support decision-making, but we argue that sufficiently advanced LLMs—such as DeepSeek V3, DeepSeek R1, and ChatGPT 4o—should be capable of playing the Werewolf game independently, given robust reasoning capabilities. While techniques like Chain-of-Thought (CoT) prompting can further enhance an LLM’s deductive reasoning by encouraging step-by-step thinking, they also introduce additional latency. Therefore, minimizing reliance on both external modules and verbose prompting strategies is essential for deploying responsive, real-time game systems where interaction speed is critical.
With the continuously improving reasoning capabilities of large language models (LLMs), recent studies have demonstrated that fine-tuned or state-of-the-art LLMs can achieve competitive performance in poker game scenarios [4, 5]. We anticipate that, as LLMs become more proficient in reasoning and in-context learning, fine-tuning and advanced prompting techniques will become increasingly unnecessary across many strategic settings. Motivated by this trend, we propose a straightforward Werewolf game framework in which we provide each simulated player with relevant context, such as game rules and general strategies, directly through prompts. The LLMs are then tasked with role-playing distinct player personas, strategically interacting to create an engaging and dynamic game experience.
In addition, the verbal communication aspect of the Werewolf game as an essential feature of its social dynamics has been largely overlooked in existing frameworks [2, 3]. To address this, we fine-tuned and integrated a text-to-speech (TTS) module into our system, enabling gameplay in both textual and auditory formats. Importantly, while audio output enhances immersion, all core reasoning remains text-based.
2 Frameworks
Our proposed Werewolf game system is composed of the following modules:
Werewolf Game: The game pipeline that manages the game flow, executes player’s actions, and judge the game status.
Players: The human or LLM role-played game players, agents with varying action space depending on their assigned roles.
User Engagement: A pipeline that parallels LLM and TTS generations.
These modules operate in parallel to create a coherent, vocal and engaging Werewolf game experience, enabling smooth interaction between game agents and users while maintaining the natural dynamics of a social deduction game.
2.1 Werewolf Game
2.1.1 Players
We define human and AI players, where AI players are role-played by LLM. AI players’ names are prepared in order to demonstrate the TTS avatars we tuned, including Jay Chow, Jack Ma, Ma Baoguo, Yang Mi, Luo Xiang, Shen Yi, and Donald Trump. The prepared names are either famous Chinese businessman, actors, influencers, or famous politician. This design choice heavily relies on the outcome of TTS module. Meanwhile, human players have the automacy to enter their preferred names in the game. The game requires at least players to play.
2.1.2 Roles
Only roles are available for the demonstration purposes since the complexities continue exploding when introducing more roles. In this simplest Werewolf game setup, the action space for each roles are notated as:
Villager:
Werewolf:
Seer:
Witch:
Here, vote refers to the act of selecting a player’s name for elimination after the discussion phase of the game. Reason denotes the verbal exchange during which players debate and identify potentially suspicious individuals, typically those suspected of being werewolves, thereby supporting their voting decisions. Reveal represents the special ability of the Seer, who, during the night phase, can uncover the hidden role of one player per round. Intuitively, kill indicates the target selected by the assigned Werewolves for elimination during the night phase of each round.
We further define two critical resources associated with the Witch’s actions: , denoting the number of remaining cures available to save a player targeted by the Werewolves, and , indicating the number of poisons that can be used to eliminate a player during the night. The Witch’s night action requires two sequential decisions formatted as: Cure/None, NameToEliminate/None.
During the day phase, active players are required to reason collaboratively through discussion, followed by a voting phase in which they vote to eliminate the most suspicious player based on inconsistencies or contradictions observed during the conversation. In the night phase, active players with special roles may perform their designated actions in a predefined sequence.
2.2 Game Flow
We encapsulate a single gameplay of the Werewolf game into a class, where player roles are automatically assigned and essential components, such as the game log, are initialized by default. Within this class, the game’s continuation status is managed by the game.judge function, which determines whether the game should proceed or end based on whether the number of active werewolves significantly exceeds that of the remaining active players. The framework evaluates the game status twice per full cycle (i.e., one night phase and one day phase): once immediately after the night phase, and again after the day phase. During the night phase, actions are executed in a predefined order: werewolves (kill) → seer (reveal) → witch (night). In the day phase, the discussion stage comes first, where all active players reason about who the werewolves might be based on prior interactions. This is followed by a voting stage where each player votes to eliminate the player they find most suspicious. We summarize the streamlined game flow as follows.
2.3 LLM Role-play
Unlike traditional agentic system, LLM-powered agent systems excels at various aspects from dynamically following and executing the instructions, to interact with the environment that requires solid reasoning capabilities [6]. The capabilities of such LLM agents largely rely on the the capabilities of LLMs the system integrated [2]. We integrated one of the most state-of-the-art LLMs DeepSeek V3 compromising longer response time of DeepSeek R1 which is designed for reasoning-intensive tasks, and higher costs of OpenAI models. However, we still integrated the portals for other LLMs from HuggingFace, llama cpp, and OpenAI for future experiment and development.
2.3.1 Variables
We list the below variables for LLMs to play the Werewolf game for solid reasoning, and only when they are necessary to their assigned roles:
active_players: Players who are currently active (i.e., active status in the game).
werewolves: List of players who are assigned with Werewolf roles.
seer_dict: The seer’s privately revealed information about other players’ roles, only visible to the seer.
werewolf_log: The action history of the werewolves(e.g. what did they do previously in the night phase), visible only to the werewolves.
witch_log: The action history of the witch, visible only to the witch.
log: The public game history, including eliminations, votes, and discussions, visible to all players.
2.3.2 Prompt Design
We assume the LLMs are as capable to play the Werewolf game without additional complex design such as Chain-of-Thoughts(CoT) prompting or ReAct prompting, considering DeepSeek V3 is distilled from DeepSeek R1[7]. We simply prompt the LLM by providing the information of players to be role-played, description of the action space, and guidance of the reply format, and very little guidance about the game rules. Below is one example of the prompt design for discussion phase.
2.4 Text-to-Speech Module
The Text-to-Speech (TTS) module implemented in this framework leverages the GPT-SoVITS[8], a state-of-the-art model combining Generative Pre-trained Transformer (GPT) techniques with Speaker-adaptive voice conversion (SoVITS) technology. This module enables high-quality voice synthesis and cloning from minimal audio input only in one shot, significantly reducing the challenge of collecting training data.
Key features of the implemented TTS module include:
Rapid Training: Only approximately one minute of speech data is required to achieve effective fine-tuning, allowing rapid deployment.
Zero-shot Capability: With as little as five seconds of speaker audio, the module can perform robust text-to-speech conversion, enabling quick adaptation to new voice profiles.
Cross-Language Support: The module supports multilingual synthesis, accommodating Chinese, English, Japanese, and other languages, enhancing versatility for diverse linguistic applications. Currently we have only set up English and Chinese modes and fine-tuned the voice packs targeting Jay Chow, Jack Ma, Ma Baoguo, Yang Mi, Luo Xiang, Shen Yi, and Donald Trump.
The TTS module was fine-tuned specifically for our framework to (1) optimize voice quality and (2) bring familiar voices into the game, enhancing the overall player experience. The streamlined training and inference pipeline significantly improves both efficiency and flexibility, enabling targeted and customizable voice synthesis for a variety of use cases.
2.5 User Engagement
2.5.1 Latency
We take into account that the process of TTS generating speech takes a certain amount of time. Instead of using a completely linear process (LLM generates a piece of text and then enters the TTS function for audio output, and then proceeds to the next stage of LLM text generation after all the audio output is completed), we processed LLM and TTS generation in parallel. The LLM generation operates in flush mode, where tokens are produced incrementally rather than waiting for the entire output to be generated. The system continuously listens to the generated text and triggers the TTS model to produce audio output either at clause boundaries, identified by commas, or at full sentence boundaries, identified by periods. The resulting audio segments are managed by a queue that ensures they are played sequentially, maintaining the natural flow of speech.
The advantage of this is that when presenting the game in audio format, this processing pipeline significantly reduces perceived response latency by allowing speech to begin as soon as the first tokens are generated, rather than waiting for the full output. This enables more natural and immersive interactions, closely mimicking human conversation patterns which is crucial in a social deduction setting where timing and tone matter. By chunking speech at clause or sentence boundaries, the system enables more natural and fluid dialogue. The audio queue allows smoother turn-taking, potential interruptions, and scalability in multi-agent settings. This design also helps keep computational resources efficiently utilized.
2.5.2 Scalability
We set aside a parameter to select the number of human players, which brings higher scalability to the program, and allows us to select one or more human users to play the game with other AI players, which also fits the real-life scenario where the user wants to play the game by himself or with 1-2 friends, but the number of people is too limited to carry out the full game flow.
We have also kept the prompt and rules definitions for the various game characters separately, which also creates the possibility of adding more game characters (such as hunters, guards, wolf lords, etc.) at a later date.
3 Qualitative Results
Overall Latency
During qualitative testing, we observed minor latency only at the beginning of the discussion phase, typically in the first few sentences spoken by the first player to respond. This slight delay stems from the initial overhead of triggering TTS generation and starting audio playback. However, once initialized, the system runs smoothly, as the audio playback duration generally exceeds the time required for both LLM generation and TTS synthesis. This allows subsequent outputs to be processed in parallel and queued efficiently, resulting in consistently low-latency, natural interactions for the remainder of the game flow.
AI Players’ Strategies and Performance
Across qualitative evaluations, our Werewolf game framework reveals that advanced LLM role-play exhibit strong autonomous reasoning and strategic gameplay, even without reliance on external tools. These agents demonstrate role-consistent behaviors, such as cautious information disclosure by Seers, resource-aware decision-making by Witches, and persuasive deception by Werewolves, indicating an emergent understanding of social deduction dynamics.
Notably, these models often take initiative in guiding discussions, shift suspicion based on evolving context, and modulate their language style to influence or evade others. Such behaviors suggest not only role awareness but also a degree of leadership and social alignment typically associated with experienced human players.
In contrast, when running less capable LLMs (e.g., smaller-scale or quantized versions), we observe a significant drop in coherence and strategic consistency. These models tend to produce repetitive, generic statements, fail to adapt to game state changes, and often overlook their special abilities entirely. Their discussions lack depth, and they rarely sustain deception or form persuasive arguments.
These findings reinforce that high-capacity LLMs can serve as autonomous, strategic agents in multi-role, socially interactive environments, demonstrating emergent leadership, coordination, and deception purely through language modeling.
API Caching
During our qualitative benchmarks, we observed that interrupting the game flow, specifically during the first night phase and immediately starting a new game could result in the API returning text generated for the previous session. This caused inconsistencies with the current game’s context. Initially, we suspected the cause is from improper handling of game logs or prompt construction. However, targeted testing with smaller local LLMs (e.g., QWen 1.5B Quant 8) failed to reproduce the issue, suggesting the problem was not on the framework side. We ultimately attributed this behavior to the API server, which, although not documented, appears to compare newly queried prompts with prior ones and return cached generations if the similarity is deemed sufficiently high.
AI Hallucination
We also observed hallucinations in the more recent version of DeepSeek-V3-0324, where the model role-played game characters by associating them with real-world public figures and made assumptions based on those identities. For instance, the model might consider Jack Ma suspicious solely because he is a prominent businessman, or deem Donald Trump untrustworthy due to his political background, while portraying Jay Chou as less suspicious because of his reputation in the music industry. This behavior persisted even after we carefully tuned the prompts to mitigate such biases. However, the issue occurred less frequently when using an earlier version of the model (DeepSeek-V3-1224) and other LLMs such as ChatGPT 4o even when without the specific prompt tuning.
4 Limitations
While our system demonstrates promising autonomous behavior in LLM-driven Werewolf gameplay, several limitations remain. First, it is not a fully agentic architecture, since our design does not incorporate external tools such as retrieval-augmented generation (RAG), persistent memory, or structured knowledge bases. All reasoning is conducted via in-context learning, which makes the system inherently constrained by the context length limitations of different LLMs. As the game progresses, earlier dialogue and decisions may be truncated, potentially impacting consistency, recall, and long-term strategic reasoning.
Second, the current implementation supports only a subset of standard Werewolf roles (e.g., Seer, Witch, Werewolf, Villager). Expanding the system to include more complex roles (e.g., Hunter, Cupid, Guard) would introduce deeper interdependencies and increase the difficulty of role alignment, offering a stronger test of agent adaptability.
Third, if our quantitative evaluations primarily rely on win/loss outcomes, such metric can introduce bias. A team victory may reflect the opposing side’s failure rather than superior strategic behavior. Future evaluations should incorporate finer-grained behavioral metrics such as reasoning depth, influence in discussion, and success in deception or alignment as a valid and comprehensive quantitative evaluations.
Fourth, while the system supports audio output via TTS for immersive interaction, it currently lacks voice input. This limits its usability in real-time spoken interactions and reduces the naturalness of gameplay in multiplayer or human-in-the-loop settings.
5 Conclusion
Our framework demonstrates that large language models with strong reasoning capabilities can autonomously and effectively participate in the Werewolf game, exhibiting strategic thinking, role-based adaptation, and socially coherent dialogue without the need for external modules or structured planning systems. By integrating a fine-tuned text-to-speech (TTS) module and leveraging a parallel processing design, the system delivers low-latency, natural audio output that enhances immersion and real-time interaction. These components together create a compelling and engaging game experience, highlighting the potential of modern LLMs not only as conversational agents but as active participants in complex, multi-agent social environments.